# Denominators in Lusztig’s asymptotic Hecke algebra via the Plancherel formula

### v5

@@ 3.4 Proof of Theorem 13 for $\mathrm{GL}_{n}$
Let $\tilde{W}$ be of type $\tilde{A}_{n}$. Then statements 1 and 2 in the statement of Theorem 2 are true, together with a stronger version of statement 3: Let $u=(r_{1},\dots,r_{k})$ be a unipotent conjugacy class in $\mathrm{GL}_{n}(\mathbb{C})$. Let $d$ be a distinguished involution in the two-sided cell $\mathbf{c}(u)\subset\tilde{W}(\mathrm{GL}_{n})$ corresponding to $u$. Let $\pi$ be the unique family of parabolic inductions that $t_{d}$ does not annihilate. Then $\operatorname{rank}\pi(t_{d})=1$.
@@
By Corollary 2 and Propositions 2 and 3, $a_{1,x}$ is a rational function of $\mathbf{q}$ for all $w$. Then equation (27) implies that $a_{x,w}$, being a sum of rational functions with Laurent polynomial coefficients, is a rational function of $\mathbf{q}$ for all $x$. The same equation, together with the fact that $J_{\mathbf{c}}$ is a two-sided ideal for each cell $\mathbf{c}$ shows that the denominator of $a_{x,w}$ depends only on the two-sided cell containing $w$. This proves the first claim. The second claim now follows from the first claim and the first statement of Corollary 2 and the fact that $\tilde{W}$ has finitely-many two-sided cells.
@@ 3.5 Proof of Theorem 2
in $J/[J,J]\otimes\mathcal{A}$. Considering traces and invoking Proposition 3, we see that the denominator of every $a_{x,w}$ divides a power of $P_{W}(\mathbf{q})$. Therefore there is $N=N_{\tilde{W}}$ depending only $\tilde{W}$ such that we can take $P^{1}_{\mathbf{G}}(\mathbf{q})=P_{W}(\mathbf{q})^{N}$. This proves the second statement of the Theorem. The third statement was proven in Corollary 1. ∎
@@ 4 Representations with fixed vectors under parahoric subgroups
Let $\mathcal{P}$ be a parahoric subgroup of $G$. Let $\pi$ be a simple tempered representation of $G$ with $I$-fixed vectors with Kazhdan-Lusztig parameter $(u,s,\rho)$. Let $w_{\mathcal{P}}$ be the longest element in the parabolic subgroup of $\tilde{W}$ defined by $\mathcal{P}$ and $\mathcal{B}_{u}^{\vee}$ be the Springer fibre for $u$.
@@
$\left(\phi\circ()^{\dagger}\right)(C^{\prime}_{w_{\mathcal{P}}})t_{w_{P}}=(-1)^{\ell(w_{\mathcal{P}})}\sum_{\overset{d\in\mathcal{D}}{z\sim_{L}d}}h_{w_{\mathcal{P}},d,z}t_{z}t_{w_{\mathcal{P}}}=(-1)^{\ell(w_{\mathcal{P}})}\sum_{z}h_{w_{\mathcal{P}},w_{\mathcal{P}},z}t_{z}=\operatorname{vol}(\mathcal{P})t_{w_{P}},$
@@
There are five conjugacy classes of parahoric subgroup, each obtained by projection from $\mathrm{GSp}_{4}(F)$: the maximal parahoric $\mathrm{PGSp}_{4}(\mathcal{O})$, the image of the paramodular group $\mathsf{K}$ corresponding to $\{0,2\}$, the image of the Siegel parahoric subgroup $\mathsf{Si}$ corresponding to $\{1\}$, the image of the Klingen parahoric $\mathsf{Kl}$ corresponding to $\{0\}$, and the Iwahori subgroup.
@@
$t_{w_{\mathcal{P}}}$\  $\pi=\operatorname{Ind}_{P}^{G}(-)$ $a$ $\nu$ $\nu\otimes\pm\mathrm{St}_{\mathrm{SO}_{3}}$ $\tau_{\mathrm{triv}}$ $\tau_{\mathrm{sgn}}$$\pm\tau_{2}$$\pm\mathrm{St}$ $\mathcal{P}$
@@
$t_{1}$ 0 8 4 3 1 4 1 $I$

### v6

@@ 3.4 The case of $\mathrm{GL}_{n}$
Let $W_{\mathrm{aff}}$ be of type $\tilde{A}_{n}$. Then statements 1 and 2 in the statement of Theorem 2 are true, together with a stronger version of statement 3: Let $u=(r_{1},\dots,r_{k})$ be a unipotent conjugacy class in $\mathrm{GL}_{n}(\mathbb{C})$. Let $d$ be a distinguished involution in the two-sided cell $\mathbf{c}(u)\subset W_{\mathrm{aff}}(\mathrm{GL}_{n})$ corresponding to $u$. Let $\pi$ be the unique family of parabolic inductions that $t_{d}$ does not annihilate. Then $\operatorname{rank}\pi(t_{d})=1$.
@@
By Corollary 3 and Propositions 2 and 3, $a_{1,x}$ is a rational function of $\mathbf{q}$ for all $w$. Then equation (29) implies that $a_{x,w}$, being a sum of rational functions with Laurent polynomial coefficients, is a rational function of $\mathbf{q}$ for all $x$. The same equation, together with the fact that $J_{\mathbf{c}}$ is a two-sided ideal for each cell $\mathbf{c}$ shows that the denominator of $a_{x,w}$ depends only on the two-sided cell containing $w$. This proves the first claim. The second claim now follows from the first claim and the first statement of Corollary 3 and the fact that $W_{\mathrm{aff}}$ has finitely-many two-sided cells.
@@ 3.5 Proof of Theorem 2
in $J/[J,J]\otimes\mathcal{A}$. Considering traces and invoking Proposition 3, we see that the denominator of every $a_{x,w}$ divides a power of $P_{W}(\mathbf{q})$. Therefore there is $N=N_{W_{\mathrm{aff}}}$ depending only $W_{\mathrm{aff}}$ such that we can take $P^{1}_{\mathbf{G}}(\mathbf{q})=P_{W}(\mathbf{q})^{N}$. This proves the second statement of the Theorem. The third statement was proven in Corollary 2. ∎
@@ 4 Representations with fixed vectors under parahoric subgroups
Let $\mathcal{P}$ be a parahoric subgroup of $G$. Let $\pi$ be a simple tempered representation of $G$ with $I$-fixed vectors with Kazhdan-Lusztig parameter $(u,s,\rho)$. Let $w_{\mathcal{P}}$ be the longest element in the parabolic subgroup of $W_{\mathrm{aff}}$ defined by $\mathcal{P}$ and $\mathcal{B}_{u}^{\vee}$ be the Springer fibre for $u$.
@@
$\left(\phi\circ^{\dagger}(-)\right)(C^{\prime}_{w_{\mathcal{P}}})t_{w_{P}}=(-1)^{\ell(w_{\mathcal{P}})}\sum_{\overset{d\in\mathcal{D}}{z\sim_{L}d}}h_{w_{\mathcal{P}},d,z}t_{z}t_{w_{\mathcal{P}}}=(-1)^{\ell(w_{\mathcal{P}})}\sum_{z}h_{w_{\mathcal{P}},w_{\mathcal{P}},z}t_{z}=\operatorname{vol}(\mathcal{P})t_{w_{P}},$
@@
There are five conjugacy classes of parahoric subgroups, each obtained by projection from $\mathrm{GSp}_{4}(F)$: the maximal parahoric $\mathrm{PGSp}_{4}(\mathcal{O})$, the image of the paramodular group $\mathsf{K}$ corresponding to $\{0,2\}$, the image of the Siegel parahoric subgroup $\mathsf{Si}$ corresponding to $\{1\}$, the image of the Klingen parahoric $\mathsf{Kl}$ corresponding to $\{0\}$, and the Iwahori subgroup.
@@
$t_{w_{\mathcal{P}}}\$ $\pi=\operatorname{Ind}_{P}^{G}(-)$ $a$ $\nu$ $\nu\otimes\pm\mathrm{St}_{\mathrm{SO}_{3}}$ $\tau_{\mathrm{triv}}$ $\tau_{\mathrm{sgn}}$    $\pm\tau_{2}$ $\pm\mathrm{St}$ $\mathcal{P}$
@@
$t_{1}$ 0 8 4 3 1 4 1 II
Acknowledgements.
The author thanks Alexander Braverman for introducing him to this material and the problem, for suggesting the use of the Plancherel formula, and for helpful conversations. The author also thanks Roman Bezrukavnikov, Itai Bar-Natan, Gurbir Dhillon, Malors Espinosa-Lara, Julia Gordon, George Lusztig, and Matthew Sunohara for helpful conversations. The author thanks Maarten Solleveld for explaining the generalization in [Solleveld] of the results of [fomFormalDegUnip], suggesting the reference [APM], and for careful reading of an early version of this manuscript. The author owes a great debt to the anonymous referee for careful reading resulting in many corrections and improvements both mathematical and expositional.
Competing interests:
The author declares none.